OpenStudy (anonymous):

integral of [x (tan^2 x)] dx

OpenStudy (anonymous):

\[\int\limits_{0}^{\Pi/3} x \tan^2x dx\] that's the question.

OpenStudy (anonymous):

try a trig identity (secx)^2-(tanx)^2 =1 so \[\int\limits_{0}^{\pi/3} x(\sec^{2}x-1)dx\]

OpenStudy (anonymous):

And??? How to continue with that?

OpenStudy (anonymous):

break up the integral and you could try integration by parts

OpenStudy (anonymous):

thanks.

OpenStudy (anonymous):

\[\sec^{2} x = \frac{1}{cos^{2}x} \] Also \[cos^{2}x = \frac{1+ cos 2x}{2}\] and further break the integral up